mary has 3.00 in nickels, dimes, and quarters. If she has twice as many dimes as quarters and five more nickels than dimes, how many coins of each type does she have?
Answer by stanbon(75887)

mary has 3.00 in nickels, dimes, and quarters. If she has twice as many dimes as quarters and five more nickels than dimes, how many coins of each type does she have?
----
Equations:
Value Eq::: 5n + 10d + 25q = 300
Quantity Eq: d = 2q
Quantity Eq: n = d+5
---------------------------
Substitute for n and for d, and solve for "q":
5(d+5) + 10(2q) + 25q = 300
---
5(2q+5) + 20q + 25q = 300
---
10q+ 25 + 45q = 300
55q = 275
q = 5 (# of quarters)
----
d = 2q = 2*5 = 10 (# of dimes)
----
n = d+5 = 15 (# of nickels)
=====================================
Cheers,
Stan H.